OpenStudy (anonymous):

how is this worked out 5,000 square feet of store space for $40,000 per year what is the annual expense for 1,200 sq ft thanks in advance

OpenStudy (anonymous):

karie I lost your response

OpenStudy (anonymous):

i got it thanks, must have been using the wrong numbers on the keyboard :)

OpenStudy (nincompoop):

5000/40000 * 1200

OpenStudy (anonymous):

5000/40,000 = 1200/x solve for x

OpenStudy (anonymous):

thanks all brain freeze

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} ft^2&price\\ \hline\\ 5000&40000\\ 1200&x \end{array}\implies \cfrac{5000}{1200}=\cfrac{40000}{x}\)

OpenStudy (nincompoop):

that's interesting